How do I calculate the Operating Expenses on Cash Flow Statement?METRO MEATS LTD P&L statement for 2015 METRO MEATS LTD Statement of Financial Position as at 31 December Sales revenue 250,00
Answer #1

Original cost of equipment sold = $18,000

Carrying value of equipment = $8,500

Accumulated depreciation on equipment sold = Original cost of equipment sold - Carrying value of equipment

= 18,000 - 8,500

= $9,500

Depreciation expense for the current year = Accumulated depreciation on December 31, 2015 + Accumulated depreciation on equipment sold - Accumulated depreciation on December 31, 2014

= 20,000 + 9,500 - 24,000

= $5,500

Selling expenses = $18,000

Selling expenses include depreciation expense

Administrative expenses = $6,000

Operating expenses paid = (Selling expenses - Depreciation expense) + Administrative expenses

= 18,000 - 5,500 + 6,000

= $18,500

Please ask if you have any query related to the question. Thank you